Grade 1110 grade maths

A blacksmith has a rectangular sheet of Iron. He has to make a cylindrical vessel both circular ends are closed. When he minimizes the wastage of the sheet of iron, then what is the ratio of the wastage to the utilised area of sheet?

Ayush
10 Years agoGrade 11

3 Answers

Sourabh Singh
10 Years ago
zThe length of the square sheet will be equal to 4r + h .
From this information you can relate both the values in a common variable and you will get the ratio.
Ayush
10 Years ago
why the length will be 4r+h.............................................................................................................?
 
Akhil
7 Years ago
Let length l
 Breadth b hence radius
r=b/4
h=b=4r
2πr=l-2r
58r/7=l
lb=232r²/7
2πrh+2πr²=176r²/7 + 44+r²/7=220r²/7
Difference=12r²/7
Ratio=12/220=3/55